How to verify that the values of the variables listed are solutions of the system of equations?
{█(x+y=2@x-y= -10)┤
X = -4, y = 6
replace your variables with your values:
x+y = 2 becomes -4 + 6 = 2 which is a true statement
x-y + -10 becomes -4 (+) -6 = -10 which is also a true statement

Answer by TSloan(1)

replace your variables with your values and see if the statements are true or false:
x+y = 2 becomes -4 + 6 = 2 which is a true statement
x-y + -10 becomes -4 (+) -6 = -10 which is also a true statement